# Observation, Analysis, and Solution: Exploring Strong Lightweight Vision Transformers via Masked Image Modeling Pre-Training

### v1

@@ Abstract
Masked image modeling (MIM) pre-training for large-scale vision transformers (ViTs) in computer vision has enabled promising downstream performance on top of the learned self-supervised ViT features. In this paper, we question if the extremely simple ViTs’ fine-tuning performance with a small-scale architecture can also benefit from this pre-training paradigm, which is considerably less studied yet in contrast to the well-established lightweight architecture design methodology with sophisticated components introduced. By carefully adapting various typical MIM pre-training methods to this lightweight regime and comparing them with the contrastive learning (CL) pre-training on various downstream image classification and dense prediction tasks, we systematically observe different behaviors between MIM and CL with respect to the downstream fine-tuning data scales. Furthermore, we analyze the frozen features under linear probing evaluation and also the layer representation similarities and attention maps across the obtained models, which clearly show the inferior learning of MIM pre-training on higher layers, leading to unsatisfactory fine-tuning performance on data-insufficient downstream tasks. This finding is naturally a guide to choosing appropriate distillation strategies during pre-training to solve the above deterioration problem. Extensive experiments on various vision tasks demonstrate the effectiveness of our observation-analysis-solution flow. In particular, our pre-training with distillation on pure lightweight ViTs with vanilla/hierarchical design ($5.7M$/$6.5M$) can achieve $79.4\%$/$78.9\%$ top-1 accuracy on ImageNet-1K, which is comparable to the current SOTA networks with sophisticated architecture design. It also improves the corresponding supervised pre-training baselines to achieve $+2.1$/$+5.3$ box AP and $+2.0$/$+4.4$ mask AP gains on the COCO detection and segmentation tasks. Our pre-training also enables SOTA performance on the ADE20K semantic segmentation task ($42.8\%$ mIoU) and LaSOT visual tracking task ($66.1\%$ AUC) in the lightweight regime. The latter even surpasses all the current SOTA lightweight CPU-realtime trackers.
Index Terms:
Self-supervised learning, masked image modeling, vision transformers, lightweight networks, knowledge distillation.
@@ 1 Introduction
On one hand, we surprisingly find that if proper pre-training is adopted, even the extremely simple lightweight ViTs with vanilla design show comparable performance to the current SOTA ViT derivatives with delicate design on ImageNet. Without bells and whistles, our implementation with MIM pre-training can achieve $78.0\%$ top-1 accuracy (see Tab. I) by using a slightly modified version of [16] with vanilla lightweight architecture, namely ViT-Tiny (5.7M) in this paper. This finding is intriguing since the results indicate that proper pre-training can bridge the performance gap between vanilla ViT architectures and delicately designed ones in the lightweight regime to a great extent. Although the latter ones may also enjoy high accuracy and attractive floating point operation (FLOP) counts, some of them (e.g., EdgeViT-XS [23], MobileViT-S [21], Mobile-Former-294M [22], SeaFormer-L [27]) with added complexity from specialized modules make the inference speed slower overall (higher latency, see Tab. VII). This interesting finding, however, does not match the practice of lightweight ViT pre-training in [33], which claims that MIM pre-training can even hurt lightweight ViT fine-tuning accuracy on ImageNet. On the other hand, we systematically observe different behaviors between MIM and CL pre-training with respect to the downstream fine-tuning data scales, i.e., MIM pre-training generally outperforms CL pre-training on the data-sufficient ImageNet classification task, while showing significantly worse performance on the data-insufficient downstream tasks.
@@
Based on the above analyses, we have developed a decoupled distillation strategy for better MIM pre-training of lightweight ViTs, which not only learns semantics at an abstract level relevant to recognition in higher layers (see Tab. VI for the significantly improved performance on data-insufficient downstream classification tasks and dense prediction tasks), but also makes the downstream models preserve the above useful locality inductive bias (see Tab. VI for the improved performance on the data-sufficient ImageNet classification task). We also experimentally demonstrate the effectiveness of our observation-analysis-solution flow by exploring a simple, efficient, lightweight, yet accurate hierarchical ViT from Hiera [14], namely Hiera-Tiny (6.5M). In particular, our pre-training with distillation on the pure ViT-Tiny and Hiera-Tiny can achieve $79.4\%$/$78.9\%$ top-1 accuracy on ImageNet-1K, which is comparable to the current SOTA networks with sophisticated architecture design, including some high-performing lightweight ConvNets. It also improves the corresponding supervised pre-training baselines to achieve $+2.1$/$+5.3$ box AP and $+2.0$/$+4.4$ mask AP gains on the COCO detection and segmentation tasks. Our pre-training also enables SOTA performance on the ADE20K semantic segmentation task ($42.8\%$ mIoU) and LaSOT visual tracking task ($66.1\%$ AUC) in the lightweight regime. The latter even surpasses all the current SOTA lightweight CPU-realtime trackers. Our improved code and raw results are released at https://github.com/wangsr126/mae-lite.
@@ 2.2 Vision Transformers
Vision transformers (ViTs) [15] is the pioneering work that applies a transformer architecture (a stack of attention modules [68]) on image patches and shows very competitive results by introducing mean color as prediction targets. Without using any convolution to have an inductive bias [^1] Different inductive biases can be injected into different learning algorithms through architectural choices, the objective function, the curriculum, or the optimization regime. [69] to focus on spatially neighbor image elements/tokens, this preliminary attempt requires a large and curated dataset to regularize the training and make the learned transformer effective. Recently, the performance of ViTs has been largely improved thanks to better training recipes [16, 70, 71]. For instance, DeiT in [16] introduces a ConvNet teacher to help the supervised training of the ViT student with vanilla architecture to acquire some inductive biases. Besides, better architecture design is explored in some ViT derivatives [18, 20] which focus on improving self-attention mechanism to boost performance. For instance, CaiT [20] introduces the LayerScale and class-attention designs for going deeper with ViTs while XCiT [18] replaces the self attention with a novel cross-covariance attention for reducing the quadratic complexity of self attention to linear complexity.
@@
Lightweight models are friendly to resource-constrained devices, so most of the above ViT derivatives have either provided their corresponding lightweight versions or been specifically designed for a better efficiency-accuracy trade-off. For example, the lightweight regime of Mobile-Former [22] covers the computational cost range from 26 MFLOPs to 508 MFLOPs. For LeViT [17], it covers the parameter number range from 7.8 M to 39.1 M. Although Mobile-Former has relatively fewer network parameters and costs significantly fewer FLOPs than LeViT, the latter enjoys much higher inference speed thanks to the ConvNet’s clothing design without using too many complicated operators. By getting rid of the complex trade-offs among model size (or parameter number), FLOPs, inference speed (or latency), and accuracy of the above ViT derivatives with delicate design, a preliminary version of this paper published in ICML’23 [98] is the first to systematically study the effects of applying SSL pre-training on lightweight ViTs with vanilla architecture design. It is demonstrated that proper pre-training based on the proposed distillation strategy can even close the performance gap between vanilla ViT architectures and delicately designed ones in the lightweight regime to a great extent. The present work goes further to investigate more typical MIM pre-training methods to provide more comprehensive analyses, propose an improved decoupled distillation strategy to unleash the great potential of pre-training, and provide more experimental results to make the conclusion more convincing. The new results especially on the downstream semantic segmentation and visual tracking tasks and using another extremely simple lightweight hierarchical ViT as the experimental unit are also impressive.
@@ 2.3 Knowledge Distillation
Knowledge distillation (KD), a mainstream approach for model compression [99, 100, 101], is best known as an effective method for taking advantage of a large number of parameters in a teacher model during training, while having an optimized smaller and efficient student model during inference. Beyond model compression, KD is also reckoned to be able to combine strengths of different learning algorithms [69] by transferring the effect of inductive biases between different models. It is claimed that having the right inductive biases is essential for obtaining high performance when data or compute is a limiting factor. For instance, the distillation-based supervised training (or fine-tuning) in DeiT [16] achieves better accuracy on vanilla ViTs by adopting a ConvNet as the teacher. Extensive methodologies have also been proposed to improve the distilling effectiveness. In [100], the mimic error of a small student network towards the soft logit output of a larger teacher network is adopted to guide the transferring of the knowledge. FitNet [101] proposes to distill semantic information from intermediate features, and a mean-square-error-based hint loss is adopted. Besides, the channel attention [102] and the relation of examples [103] are considered as the learned knowledge to be transferred. Moreover, KD is widely adopted in various tasks apart from classification tasks, e.g., object detection [104] and semantic segmentation [105].
@@
In this paper, we focus on further improving the performance of MIM pre-trained lightweight ViTs by distillation, which is developed following the guidance of our thorough analyses. Our MIM pre-training with distillation not only helps the pre-trained lightweight ViTs to learn semantics at an abstract level relevant to recognition in higher layers, but also makes the downstream models preserve the useful locality inductive bias benefiting from MIM pre-training. Although our distillation bears some similarities to the recent TinyMIM work [33], the whole of our observation-analysis-solution flow is orthogonal to it as it aims to systematically study the distillation framework for MIM pre-trained ViTs.
@@ 3 Observation
In this section, we carefully adapt various typical pre-training methods to the lightweight regime of pre-training ViTs, benchmark the pre-trained models with a small-scale architecture on ImageNet, and further evaluate their transferability to other datasets and tasks, which addresses a little-explored question: how well does pre-training work on lightweight ViTs with vanilla architecture design, especially for MIM pre-training?
@@ 3.1 Preliminaries and Experimental Study Design
Compared Methods. We develop and carefully adapt various recently popular MIM pre-training methods (mainly for the large models) to apply them on ViT-Tiny, i.e., the MAE [12] and SimMIM [28] methods which aim to predict raw RGB pixels of masked patches, and some methods with other pre-defined targets for prediction (e.g., discrete tokens in BEiT [11], features from a momentum teacher in BootMAE [2], and HoG features in MaskFeat [72]). We also benchmark them along with two typical CL pre-training methods (i.e., MoCo-v3 [6] and DINO [1]) and the fully-supervised pre-training methods as the baselines.
@@
Baseline without pre-training. Largely following the recipe in [16], we train a ViT-Tiny model from scratch for 300 epochs on the training set of ImageNet-1K (dubbed IN1K), which serves as the baseline without pre-training in our study. It achieves 74.5% top-1 accuracy on the validation set of IN1K, surpassing that achieved based on the original architecture with 3 heads (i.e., 72.2% in [16]). It further reaches 75.8% by adopting our improved training recipe with some optimized hyper-parameters for augmentations (i.e., weaker regularization and augmentation than training large-scale models). Besides, layer-wise lr decay is also taken into consideration during fine-tuning following [11, 12]. Please refer to Sec. A.1 for more details. This can be seen as a strong baseline to examine the pre-training in our study because it even surpasses the original transformer-specific distillation-based version in [16] (i.e., 74.5%).
@@
TABLE I: Comparison Across Different Pre-Training Methods by Evaluating the Pre-Trained Models on Data-Sufficient IN1K. We Report Top-1 Accuracy on Its Validation Set. Our Modified ViT-Tiny with Vanilla Architecture Is Used As the Experimental Unit. The Pre-Training Time Is Measured on an 8$\times$×V100 GPU Machine. ‘Ori.’ Represents the Original Training/Fine-Tuning Recipe in DeiT [16] and ‘Ours’ Represents Our Improved Recipe Over It (See Sec. A.1).
Models Pre-Training Fine-Tuning Top-1 Acc.
@@
Vanilla Architecture CL Self-Supervised
DINO [1] IN1K w/o labels 400 83 Ours 300 77.2
@@
CL pre-training baselines. To adapt MoCo-v3 [6] and DINO [1] to ViT-Tiny, we re-implement them with ViT-Tiny as encoder and largely follow the original setups in their papers. We conduct 400-epoch pre-training on IN1K for them. In [6], it is claimed that instability is a major issue that impacts self-supervised ViT pre-training and causes mild degradation in accuracy, and a simple trick by adopting fixed random patch projection (the first layer of a ViT model) is proposed to improve the stability in practice. However, we find that instability is not the main issue for MoCo-v3 adaptation to ViT-Tiny in our study, and higher performance is achieved with a learned patch projection layer. In [1], DINO is interpreted as a form of knowledge distillation with no labels, where a centering and sharpening of the teacher output is used to avoid collapse.
MIM pre-training methods. We carefully adapt MAE [12], BEiT [11], BootMAE [31] and MaskFeat [32] to ViT-Tiny by investigating the different effects of several basic factors and components in their frameworks. For instance, as a key component, decoder determines the semantic level of the learned latent representations. It has been validated that MIM works well across a wide range of the decoder’s width and depth under the fine-tuning protocol when the encoder is large. However, we experimentally find that MIM prefers a much more lightweight decoder when the encoder is altered to ViT-Tiny, i.e., a decoder consisting of one transformer block with the head number and width set to 3 and 96. We also conduct a small sweep over 5 masking ratios $\{0.45,0.55,0.65,0.75,0.85\}$ and find that 0.75 achieves the best performance. Other setups in the above papers remain intact including the optimizer, learning rate, batch size, argumentation, etc. As for SimMIM [28] adaptation to ViT-Tiny, we rigorously follow the original settings in the paper. We also do MIM pre-training on IN1K for 400 epochs.
3.2 Fine-Tuning Evaluation on ImageNet Classification
Fine-Tuning Evaluation. We follow the practice of supervised ViT-Tiny training with our improved recipe in Sec. A.1 for fine-tuning evaluation, though the decay coefficient in the layer-wise lr decay schedule is separately optimized for different pre-trained models. Besides, we use global average pooling (GAP) after the final block during the fine-tuning of both the MIM and CL pre-trained models, which is, however, not the common practice for MoCo-v3 [6]. We adopt GAP for MoCo-v3 as it significantly helps to surpass the model using the original configuration based on a class token (76.8% vs. 73.7%) for our used lightweight ViT-Tiny.
Observation 1. As shown in Tab. I, almost all of the compared pre-training methods, including both the fully-supervised and self-supervised ones, can achieve better downstream fine-tuning performance on the ImageNet classification task than the random initialization paradigm without pre-training, whilst MIM pre-training with moderate pre-training cost consumed outperforms the compared pre-training baselines. As we scale up the supervised pre-training duration on IN21K from 30 to 300 epochs, we find that the 10×10$\times$10 × longer schedule can achieve a comparable performance (i.e., $77.8\%$) to the MIM pre-training, while the pre-training cost gets much higher due to the significantly increased pre-training time, let alone the used expensive labels. We also conduct the supervised pre-training with 300 epochs on IN1K using our fine-tuning recipe to better show the benefits of self-supervised pre-training or increasing the training data size. All the results indicate that the extremely simple lightweight ViTs with vanilla design have great potential which can be unleashed via proper pre-training.
TABLE II: Effect of Varying the SSL Pre-Training Data Size and Class Distribution Based on ViT-Tiny. We Follow the 400-Epcho Self-Supervised Pre-Train, 300-Epoch Supervised Fine-Tune Paradigm. Top-1 Accuracy on the Validation Set of IN1K Is Reported.
@@ 3.2 Fine-Tuning Evaluation on ImageNet Classification
TABLE III: Transfer Evaluation of Our Modified ViT-Tiny on Some Data-Insufficient Downstream Classification Tasks and Dense Prediction Tasks. Top-1 Accuracy Is Reported for Classification Tasks, mIoU Is for ADE20K Segmentation, AUC Is for LaSOT Tracking, and AP Is for COCO Detection (Det.) and Segmentation (Seg.).The Dataset Scale Is Represented As (Train-Size Per Epoch/#Train-Classes).

### v2

@@ Abstract
Masked image modeling (MIM) pre-training for large-scale vision transformers (ViTs) has enabled promising downstream performance on top of the learned self-supervised ViT features. In this paper, we question if the extremely simple lightweight ViTs’ fine-tuning performance can also benefit from this pre-training paradigm, which is considerably less studied yet in contrast to the well-established lightweight architecture design methodology. We use an observation-analysis-solution flow for our study. We first systematically observe different behaviors among the evaluated pre-training methods with respect to the downstream fine-tuning data scales. Furthermore, we analyze the layer representation similarities and attention maps across the obtained models, which clearly show the inferior learning of MIM pre-training on higher layers, leading to unsatisfactory transfer performance on data-insufficient downstream tasks. This finding is naturally a guide to designing our distillation strategies during pre-training to solve the above deterioration problem. Extensive experiments have demonstrated the effectiveness of our approach. Our pre-training with distillation on pure lightweight ViTs with vanilla/hierarchical design ($5.7M$/$6.5M$) can achieve $79.4\%$/$78.9\%$ top-1 accuracy on ImageNet-1K. It also enables SOTA performance on the ADE20K segmentation task ($42.8\%$ mIoU) and LaSOT tracking task ($66.1\%$ AUC) in the lightweight regime. The latter even surpasses all the current SOTA lightweight CPU-realtime trackers.
keywords:
Masked image modeling, vision transformers, lightweight networks, knowledge distillation, visual tracking
@@ 1 Introduction
On one hand, we surprisingly find that if proper pre-training is adopted, even the extremely simple lightweight ViTs with vanilla design show comparable performance to the current SOTA ViT derivatives with delicate design on ImageNet. Without bells and whistles, our implementation with MIM pre-training can achieve $78.0\%$ top-1 accuracy (see Tab. 1) by using a slightly modified version of [16] with vanilla lightweight architecture, namely ViT-Tiny (5.7M) in this paper. This finding is intriguing since the results indicate that proper pre-training can bridge the performance gap between vanilla ViT architectures and delicately designed ones in the lightweight regime to a great extent. Although the latter ones may also enjoy high accuracy and attractive floating point operation (FLOP) counts, some of them (e.g., EdgeViT-XS [23], MobileViT-S [21], Mobile-Former-294M [22], SeaFormer-L [27]) with added complexity from specialized modules make the inference speed slower overall (higher latency, see Tab. 7). This interesting finding, however, does not match the practice of lightweight ViT pre-training in [33], which claims that MIM pre-training can even hurt lightweight ViT fine-tuning accuracy on ImageNet. On the other hand, we systematically observe different behaviors between MIM and CL pre-training with respect to the downstream fine-tuning data scales, i.e., MIM pre-training generally outperforms CL pre-training on the data-sufficient ImageNet classification task, while showing significantly worse performance on the data-insufficient downstream tasks.
@@
Based on the above analyses, we have developed a decoupled distillation strategy for better MIM pre-training of lightweight ViTs, which not only learns semantics at an abstract level relevant to recognition in higher layers (see Tab. 6 for the significantly improved performance on data-insufficient downstream classification tasks and dense prediction tasks), but also makes the downstream models preserve the above useful locality inductive bias (see Tab. 6 for the improved performance on the data-sufficient ImageNet classification task). We also experimentally demonstrate the effectiveness of our observation-analysis-solution flow by exploring a simple, efficient, lightweight, yet accurate hierarchical ViT from Hiera [14], namely Hiera-Tiny (6.5M). In particular, our pre-training with distillation on the pure ViT-Tiny and Hiera-Tiny can achieve $79.4\%$/$78.9\%$ top-1 accuracy on ImageNet-1K, which is comparable to the current SOTA networks with sophisticated architecture design, including some high-performing lightweight ConvNets. It also improves the corresponding supervised pre-training baselines to achieve $+2.1$/$+5.3$ box AP and $+2.0$/$+4.4$ mask AP gains on the COCO [34] detection and segmentation tasks. Our pre-training also enables SOTA performance on the ADE20K [35] semantic segmentation task ($42.8\%$ mIoU) and LaSOT [36] visual tracking task ($66.1\%$ AUC) in the lightweight regime. In specific, our pre-trained ViT-Tiny not only outperforms the image classification-proficient LeViT [17], but also achieves comparable accuracy with the current SOTA segmentation-specific models (i.e., SeaFormer-L [27] and EfficientViT-B1 [26]) on ADE20K and even enhances the lightweight OSTrack [37] to outperform all the current SOTA lightweight CPU-realtime trackers with a new record of $66.1\%$ AUC on LaSOT. Our improved code and raw results are released at https://github.com/wangsr126/mae-lite.
This paper extends our conference version published in ICML’23 [38], which is the first to successfully perform MIM pre-training for lightweight ViT-Tiny (5.7M) with vanilla design, and has experimentally demonstrated that proper pre-training can bridge the performance gap between vanilla ViT architectures and delicately designed ones in the lightweight regime to a great extent. The present work goes further to make the following additional contributions in this journal version. (1) We investigate additional pre-training baselines (i.e., CL-based DINO and MIM-based BEiT, BootMAE, MaskFeat, SimMIM) in our study to make our observation-analysis-solution flow more conclusive. (2) We conduct the new experimental study on the ADE20K semantic segmentation and LaSOT visual tracking tasks (see Tabs. LABEL:tab:ade20ksota and LABEL:tab:lasotsota), which show that our pre-training enables the more versatile ViT model ViT-Tiny to be outstanding on both of these two tasks. (3) We apply our distillation strategies on the MAE pre-training of the recently proposed Hiera-Tiny (6.5M) to demonstrate the generalization capability of our observation-analysis-solution flow when extended to broader architecture designs. (4) We further improve the original distillation strategy (D-MAE) by solving the issue that the attention weights on the final layer of the obtained ViT-Tiny based on D-MAE exhibit some weird behavior patterns (see Fig. 10). The new decoupled distillation strategy D2-MAE not only consistently improves over D-MAE on all our evaluated datasets (see Tab. 6), but also enables the pure ViT-Tiny to achieve $79.4\%$ top-1 accuracy on ImageNet-1K, which is comparable to the current lightweight and efficient SOTA networks with sophisticated architecture design, e.g., LeViT-192 [17], FastViT-T12 [24], EfficientFormerV2-S1 [25], SeaFormer-L [27], EfficientViT-B1 [26], etc. (5) More competitors are added for comparison, and many new references are included and summarized in the related works. (6) All parts of the paper are re-organized and re-written to include the above contributions better.
@@ 2.2 Vision Transformers
Vision transformers (ViTs) [15] is the pioneering work that applies a transformer architecture (a stack of attention modules [73]) on image patches and shows very competitive results by introducing mean color as prediction targets. Without using any convolution to have an inductive bias [^1] Different inductive biases can be introduced through architectural choices, the objective function, the curriculum, or the optimization regime. [74] to focus on spatially neighbor image elements/tokens, this preliminary attempt requires a large and curated dataset to regularize the training and make the learned transformer effective. Recently, the performance of ViTs has been largely improved thanks to better training recipes [16, 75, 76]. For instance, DeiT in [16] introduces a ConvNet teacher to help the supervised training of the ViT student with vanilla architecture to acquire some inductive biases. Besides, better architecture design is explored in some ViT derivatives [18, 20] which focus on improving self-attention mechanism to boost performance. For instance, CaiT [20] introduces the LayerScale and class-attention designs for going deeper with ViTs while XCiT [18] replaces the self attention with a novel cross-covariance attention for reducing the quadratic complexity of self attention to linear complexity.
@@
Lightweight models are friendly to resource-constrained devices, so most of the above ViT derivatives have either provided their corresponding lightweight versions or been specifically designed for a better efficiency-accuracy trade-off. For example, the lightweight regime of Mobile-Former [22] covers the computational cost range from 26 MFLOPs to 508 MFLOPs. For LeViT [17], it covers the parameter number range from 7.8 M to 39.1 M. Although Mobile-Former has relatively fewer network parameters and costs significantly fewer FLOPs than LeViT, the latter enjoys much higher inference speed thanks to the ConvNet’s clothing design without using too many complicated operators. By getting rid of the complex trade-offs among model size (or parameter number), FLOPs, inference speed (or latency), and accuracy of the above ViT derivatives with delicate design, a preliminary version of this paper published in ICML’23 [38] is the first to systematically study the effects of applying SSL pre-training on lightweight ViTs with vanilla design. It is demonstrated that proper pre-training based on the proposed distillation strategy can even close the performance gap between vanilla ViT architectures and delicately designed ones in the lightweight regime to a great extent. The present work goes further to investigate more typical pre-training methods to provide more comprehensive analyses, propose an improved decoupled distillation strategy to unleash the great potential of pre-training, and provide more experimental results to make our study more conclusive. The new results on the downstream semantic segmentation and visual tracking tasks are also impressive.
@@ 2.3 Knowledge Distillation
Knowledge distillation (KD), a mainstream approach for model compression [103, 104, 105], is best known as an effective method for taking advantage of a large number of parameters in a teacher model during training, while having an optimized smaller and efficient student model during inference. Beyond model compression, KD is also reckoned to be able to combine strengths of different learning algorithms [74] by transferring the effect of inductive biases between different models. It is claimed that having the right inductive biases is essential for obtaining high performance when data or compute is a limiting factor. For instance, the distillation-based supervised training (or fine-tuning) in DeiT [16] achieves better accuracy on vanilla ViTs by adopting a ConvNet as the teacher. Extensive methodologies have also been proposed to improve the distilling effectiveness. In [104], the mimic error of a small student network towards the soft logit output of a larger teacher network is adopted to guide the transferring of the knowledge. FitNet [105] proposes to distill semantic information from intermediate features, and a MSE-based hint loss is adopted. Besides, the channel attention [106] and the relation of examples [107] are considered as the learned knowledge to be transferred. Moreover, KD is widely adopted in various tasks apart from classification tasks, e.g., object detection [108] and semantic segmentation [109].
@@
In this paper, we focus on further improving the performance of MIM pre-trained lightweight ViTs by distillation, which is developed following the guidance of our thorough analyses. Our MIM pre-training with distillation not only helps the pre-trained lightweight ViTs to learn semantics at an abstract level relevant to recognition in higher layers, but also makes the downstream models preserve the useful locality inductive bias benefiting from MIM pre-training.
@@ 3 Observation
In this section, we carefully adapt various typical pre-training methods to the lightweight regime of pre-training ViTs, benchmark the pre-trained lightweight models on ImageNet, and further evaluate their transferability to other datasets and tasks, which addresses a little-explored question: how well does pre-training work on lightweight ViTs with vanilla architecture design, especially for MIM pre-training?
@@ 3.1 Preliminaries and Experimental Study Design
Compared Methods. We develop and carefully adapt various recently popular MIM pre-training methods (mainly for the large models) to apply them on ViT-Tiny, i.e., the MAE [12] and SimMIM [28] methods which aim to predict raw RGB pixels of masked patches, and some methods with other pre-defined targets for prediction (e.g., discrete tokens in BEiT [11], features from a momentum teacher in BootMAE [2], and HoG features in MaskFeat [72]). We also benchmark them along with two CL pre-training methods (i.e., MoCo-v3 [6] and DINO [1]) and the fully-supervised pre-training methods as the baselines.
@@
Baseline without pre-training. Largely following the recipe in [16], we train a ViT-Tiny model from scratch for 300 epochs on the training set of ImageNet-1K (dubbed IN1K), which serves as the baseline without pre-training in our study. It achieves 74.5% top-1 accuracy on the validation set of IN1K, surpassing that achieved based on the original architecture with 3 heads (i.e., 72.2% in [16]). It further reaches 75.8% by adopting our improved training recipe with some optimized hyper-parameters for augmentations (i.e., weaker regularization and augmentation than training large-scale models). Besides, layer-wise lr decay is also taken into consideration during fine-tuning following [11, 12]. Please refer to Appendix A.1 for more details. This can be seen as a strong baseline to examine the pre-training in our study because it even surpasses the original transformer-specific distillation-based version in [16] (i.e., 74.5%).
CL pre-training baselines. To adapt MoCo-v3 [6] and DINO [1] to ViT-Tiny, we re-implement them with ViT-Tiny as encoder and largely follow the original setups in their papers. We conduct 400-epoch pre-training on IN1K for them. In [6], it is claimed that instability is a major issue that impacts self-supervised ViT pre-training and causes mild degradation in accuracy, and a simple trick by adopting fixed random patch projection (the first layer of a ViT model) is proposed to improve the stability in practice. However, we find that instability is not the main issue for MoCo-v3 adaptation to ViT-Tiny, and higher performance is achieved with a learned patch projection layer. In [1], DINO is interpreted as a form of knowledge distillation with no labels, where a centering and sharpening of the teacher output is used to avoid collapse.
MIM pre-training methods. We carefully adapt MAE [12], BEiT [11], BootMAE [31] and MaskFeat [32] to ViT-Tiny by investigating the different effects of several basic factors and components in their frameworks. For instance, as a key component, decoder determines the semantic level of the learned latent representations. Although MIM works well across a wide range of the decoder’s width and depth under the fine-tuning protocol when the encoder is large, we experimentally find that MIM prefers a much more lightweight decoder when the encoder is altered to ViT-Tiny, i.e., a decoder consisting of one transformer block with the head number and width set to 3 and 96. We also conduct a small sweep over 5 masking ratios $\{0.45,0.55,0.65,0.75,0.85\}$ and find that 0.75 achieves the best performance. Other setups in the above papers remain intact including the optimizer, learning rate, batch size, argumentation, etc. As for SimMIM [28] adaptation to ViT-Tiny, we rigorously follow the original settings in the paper. We also do MIM pre-training on IN1K for 400 epochs.
3.2 Fine-Tuning Evaluation on ImageNet Classification
Fine-Tuning Evaluation. We follow the practice of supervised ViT-Tiny training with our improved recipe in Appendix A.1 for fine-tuning evaluation, though the decay coefficient in the layer-wise lr decay schedule is separately optimized for different pre-trained models. Besides, we use global average pooling (GAP) after the final block during the fine-tuning of both the MIM and CL pre-trained models, which is, however, not the common practice for MoCo-v3 [6]. We adopt GAP for MoCo-v3 as it significantly helps to surpass the model using the original configuration based on a class token (76.8% vs. 73.7%) for our used lightweight ViT-Tiny.
@@ 3.2 Fine-Tuning Evaluation on ImageNet Classification
Table 1: Comparison across different pre-training methods by evaluating the pre-trained models on data-sufficient IN1K. We report top-1 accuracy on its validation set. Our modified ViT-Tiny with vanilla architecture is used as the experimental unit. The pre-training time is measured on an 8$\times$×V100 GPU machine. ‘Ori.’ represents the original training/fine-tuning recipe in DeiT [16] and ‘Ours’ represents our improved recipe (see Appendix A.1).
Models Pre-Training Fine-Tuning Top-1
@@
CL Self-Supervised
Vanilla Architecture DINO [1] IN1K w/o labels 400 83 Ours 300 77.2
@@
Observation 1. As shown in Tab. 1, almost all of the compared pre-training methods, including both the fully-supervised and self-supervised ones, can achieve better downstream fine-tuning performance on the ImageNet classification task than the random initialization paradigm without pre-training, whilst MIM pre-training with moderate pre-training cost consumed outperforms the compared pre-training baselines. As we scale up the supervised pre-training duration on IN21K from 30 to 300 epochs, we find that the 10×10$\times$10 × longer schedule can achieve a comparable performance (i.e., $77.8\%$) to the MIM pre-training, while the pre-training cost gets much higher due to the significantly increased pre-training time, let alone the used expensive labels. We also conduct the supervised pre-training with 300 epochs on IN1K using our fine-tuning recipe to better show the benefits of self-supervised pre-training or increasing the training data size. All the results indicate that the extremely simple lightweight ViTs with vanilla design have great potential which can be unleashed via proper pre-training.
Table 2: Effect of varying the SSL pre-training data size and class distribution based on ViT-Tiny. We follow the 400-epcho self-supervised pre-train, 300-epoch supervised fine-tune paradigm. Top-1 accuracy on the validation set of IN1K is reported.
@@
Observation 2. The above observation encourages us to further explore how the enhanced lightweight ViTs perform in comparison to the recent SOTA lightweight ViT derivatives (including both the hybrid and non-hybrid) as introduced in Tab. 7. It can be observed that the enhanced vanilla ViT-Tiny through MIM pre-training can also achieve SOTA performance on par with some superior ViT derivatives with comparable parameter number. It is noteworthy that these delicately designed architectures in the lightweight regime, despite lack of pre-training, achieve top performance on the ImageNet classification task by seeking to (1) extend the supervised fine-tuning duration to a much longer schedule (e.g., 1000 epochs in LeViT [17] and PiT [19]), (2) introduce distillation-driven training to regularize the fine-tuning (e.g., LeViT [17], XCiT [18], PiT [19], and CaiT [20], etc.), or (3) use some recent techniques like relative position bias [120, 121, 122, 123, 124] (e.g., LeViT [17] and Swin [91]), conditional positional encoding [126] (e.g., EdgeViT [23]) and LayerScale (e.g., XCiT [18]). The competing results of the enhanced ViT-Tiny demonstrate the benefits of studying MIM pre-training on extremely simple lightweight ViTs, which is orthogonal to the network architecture design methodology. Besides, extremely simple architectures are usually friendly to the real-world deployment by getting rid of the complicated modules existing in some ViT derivatives (e.g., EdgeViT-XS [23], MobileViT-S [21], Mobile-Former-294M [22], SeaFormer-L [27]).
Observation 3. We also empirically examine whether the self-supervised pre-training on lightweight ViTs can benefit from large-scale pre-training data. In Tab. 2, we study the effect of varying the pre-training data size and class distribution with two typical SSL methods (i.e., MoCo-v3 and MAE) applied on ViT-Tiny. We first consider the much larger dataset IN21K, however, few improvements are observed for both MoCo-v3 and MAE. We believe the limited capacity of extremely simple lightweight ViTs also bottlenecks the representation quality of self-supervised pre-training even with the “LLM-like” scaling of data. We further consider two subsets of IN1K containing 10% and 1% of the total examples (IN1K 10% and IN1K 1%) balanced in terms of classes [127] and one subset with long-tailed class distribution [128] (IN1K-LT). Surprisingly, MAE pre-training has marginal performance declines on these subsets, showing more robustness than MoCo-v3 in terms of the pre-training data size and class distribution.
3.3 Transfer Learning Evaluation on Downstream Tasks
We further examine the transfer learning performance of the pre-trained models in Tab. 3. We only consider the models pre-trained based on IN1K for simplicity, involving their transfer performance on some other downstream classification tasks and dense prediction tasks.
Table 3: Transfer evaluation of our modified ViT-Tiny on some data-insufficient downstream classification tasks and dense prediction tasks. Top-1 accuracy is reported for classification tasks, mIoU is for ADE20K segmentation, AUC is for LaSOT tracking, and AP is for COCO detection and segmentation (Det./Seg.).The dataset scale is represented as (Train-size per epoch/#Train-classes).